Ceelo Company purchased (at a cost of $23,290) and used 3,400 pounds of materials during May. Ceelo’s standard cost of materials per unit produced is based on 1 pounds per unit at a cost $7.00 per pound. Production in May was 3,340 units.



Compute the total, price, and quantity variances for materials. (Round answers to 0 decimal places, e.g. 125.)

Total Materials Variance $                                                           FavorableUnfavorableNeither favorable nor unfavorable
Materials Price Variance $                                                           UnfavorableFavorableNeither favorable nor unfavorable
Materials Quantity Variance
Answer #1
Answer
Answer :- actual cost per unit 6.85
Actual Quantity = $3400
Standard Quantity $3340
Standard cost per unit $7
1 Total cost variance = ($3340×7)-(3400×6.85)= $90 favorable
2 Material price variance = $3400(7-6.85) $510 favorable
3 Material quantity variance =7(3340-3400) = $ 420 adverse
please Like ( Positive rating )